# Designing Beyond Language: Sociotechnical Barriers in AI Health Technologies for Limited English Proficiency

### v1

@@ Abstract.
Limited English proficiency (LEP) patients in the U.S. face systemic barriers to healthcare beyond language and interpreter access, encompassing procedural and institutional constraints. AI advances may support communication and care through on-demand translation and visit preparation, but also risk exacerbating existing inequalities. We conducted storyboard-driven interviews with 14 patient navigators to explore how AI could shape care experiences for Spanish-speaking LEP individuals. We identified tensions around linguistic and cultural misunderstandings, privacy concerns, and opportunities and risks for AI to augment care workflows. Participants highlighted structural factors that can undermine trust in AI systems, including sensitive information disclosure, unstable technology access, and low digital literacy. While AI tools can potentially alleviate social barriers and institutional constraints, there are risks of misinformation and uprooting human camaraderie. Our findings contribute design considerations for AI that support LEP patients and care teams via rapport-building, education and language support, and minimizing disruptions to existing practices.
@@ 1. Introduction
Patient-provider communication is strongly linked to health outcomes and quality of care (Arora, 2003; Street Jr et al., 2009; Ong et al., 2000). Effective communication builds therapeutic relationships, satisfaction, trust, and long-term health outcomes (Street Jr et al., 2009; Ha and Longnecker, 2010), while poor communication contributes to errors, compromised patient safety, and inefficient use of resources (Vermeir et al., 2015; Tiwary et al., 2019). Language barriers in particular can exacerbate these challenges. In the U.S., over 25 million individuals aged 5 years and older report speaking English less than “very well” (U.S. Census Bureau, 2022). These individuals with limited English proficiency (LEP) are more likely to experience misdiagnoses, medication complications, poorer adherence to treatment plans, and decreased comprehension of diagnoses (Crane, 1997; Gandhi et al., 2000; Flores, 2006; Tiwary et al., 2019). Consequently, patient-provider communication, relationship-building, and shared decision-making break down (Paredes et al., 2018; Flores, 2006), reducing both patient and provider satisfaction (Al Shamsi et al., 2020).
@@
Digital health technologies, such as telehealth, online patient portals, and mobile health applications, can facilitate asynchronous communication and on-demand information access, which in principle could improve the quality and convenience of healthcare by enabling continuous and accessible support (Fatehi et al., 2020; Madanian et al., 2023; Ezeamii et al., 2024). However, LEP populations underutilize these resources due to persistent implementation barriers such as low digital and health literacy, unreliable Internet access, and ethical and privacy concerns (Ahmed et al., 2025; Khan et al., 2025; Madanian et al., 2023; Nouri et al., 2020; Payvandi et al., 2022; Eberly et al., 2020; Reed et al., 2020). Even technologies designed specifically to accommodate LEP individuals, such as multilingual patient portals, are often poorly integrated or have limited functionalities, thus discouraging use (Anthony et al., 2024).
@@
Emerging AI-powered health technologies have the potential to provide personalized and timely medical support and translation (Johnson et al., 2021; Genovese et al., 2024; Miller and Brown, 2018), enabling LEP patients to engage with the healthcare system more effectively. For example, on-demand AI translation can facilitate patient-provider conversations (Mehandru et al., 2023), and new large language model (LLM)-powered chatbots can summarize and simplify complex medical information (Guo et al., 2021) or help patients prepare for visits (Liu et al., 2025b; Samiee et al., 2025). However, most AI-powered interventions target digitally literate, English-speaking patients (Jin et al., 2024; Naamati-Schneider and Alt, 2024), and the models underpinning these interventions generally struggle to adapt to marginalized cultural (Santy et al., 2023) and language backgrounds (Basoah et al., 2025; Sap et al., 2019). Therefore, the rapid deployment of such AI-powered health interventions runs the risk of replicating or further exacerbating existing disparities (Ahmed et al., 2025).
@@ 2. RQ2:
To address the above RQs, we conducted 14 storyboard-centered interviews with individuals who support Spanish-speaking LEP individuals receiving care in the U.S. These individuals, whom we collectively refer to as patient navigators, assisted patients in various ways, including accompanying them to provider visits, translating, teaching digital skills, and bridging resources. Their role as mediators between patients and providers lends them insights into the broader contexts and relational dynamics that shape LEP patient experiences. Consequently, they can build on their experiences to reason about the feasibility and risks of developing AI technologies to facilitate equitable communication and care.
From these interviews, we identified key communication and systemic challenges that Spanish-speaking LEP patients face, including low literacy and privacy concerns. These challenges also highlighted possible risks and opportunities for using AI, such as reducing social barriers to care while potentially diminishing existing human support. Our findings highlight the importance of adapting to cultural and linguistic nuances, overcoming various types of literacy barriers, and respecting user privacy to foster trust. These findings inform practical guidelines for developing AI tools that support LEP populations. Our contributions are summarized as follows:
@@ 2.1. Mediating Clinical Communication with LEP Individuals
Effective healthcare requires two-way communication, where providers contribute medical expertise and patients share their needs and preferences, to collaboratively develop optimal care plans (Iroegbu et al., 2025; Williams et al., 2017). Strong interpersonal communication has been shown to influence patient outcomes as much as, if not more than, clinical instruction (Bartlett et al., 1984), leading to improved adherence to treatment plans (Street Jr et al., 2009), understanding of diagnoses (Kennedy MBA et al., 2014), and satisfaction (Ruben et al., 2020). Conversely, impeded communication is linked to more errors (Tiwary et al., 2019), worsened decision-making (Paredes et al., 2018), and lower rates of physician visits and preventive services (Jacobs et al., 2006; Woloshin et al., 1997; Fiscella et al., 2002). For LEP individuals, these challenges are compounded by linguistic (Schouten et al., 2020), cultural (Lazaro et al., 2024), intersectional (Parthab Taylor et al., 2013), and procedural (Frank, 2000) barriers. Therefore, there is a growing interest in improving health outcomes by studying communication strategies that integrate patients’ lived experiences and values (Seo et al., 2021).
Research shows that language-concordant care, whether via direct communication  (Molina and Kasper, 2019) or interpretation  (Dunlap et al., 2015; Eskes et al., 2013), leads to fewer communication errors (Karliner et al., 2007) and higher satisfaction and perceived quality of care (Diamond et al., 2019) for LEP individuals. Despite these benefits, interpretation services remain underutilized due to barriers such as limited availability, cost, time constraints, and policies (Khoong and Fernandez, 2021; Jaeger et al., 2019; Jacobs et al., 2004). Beyond linguistic access, broader sociocultural factors complicate LEP communication. Patients may self-censor due to stigma, fear of judgment, or mistrust of institutions (Levy et al., 2018; El-Krab et al., 2023). These dynamics highlight the need for cultural sensitivity, empathy, and attention to the broader contexts of patients’ lives.
@@ 2.2. AI Tools in Healthcare
Recent AI advancements have potential to support healthcare systems by improving clinical decision-making and workflows, resource allocations, medical imaging, patient monitoring, medical translations, and patient-provider relationships (Maleki Varnosfaderani and Forouzanfar, 2024; Shaheen, 2021; Barwise et al., 2024; Koski and Murphy, 2021). However, prior work has also identified key challenges and risks when implementing AI for healthcare, including data privacy, lack of transparency in decision-making, misinformation, bias mitigation, and cultural sensitivity  (Maleki Varnosfaderani and Forouzanfar, 2024; Germani et al., 2024; Hussain, 2024; Capraro et al., 2024; Chustecki, 2024).
In particular, AI has potential to support marginalized communities in receiving care by overcoming language barriers, providing personalized health information, and offering support in rural or low-resource areas (Genovese et al., 2024; Kaur and Mekled, 2025). Barwise et al. (2024) explored patient’s perception of using predictive models for interpreter prioritization and Mehandru et al. (2023) used physician-in-the-loop methods to improve reliance of machine translation in settings where no human translators were available. Other work has also explored using chatbots for pre-clinical consultations (Liu et al., 2025b; Samiee et al., 2025) and helping patients read physician notes (Kambhamettu et al., 2025), discharge instructions (Zaretsky et al., 2024), and medical information (Guo et al., 2021). While AI presents many exciting possibilities to support patient care, the majority of such interventions are focused on English or translating into English, posing issues for low-resource languages such as reinforcing biases and overlooking cultural and infrastructural challenges may cause AI to exacerbate rather than alleviate inequities in underserved populations (Wasi, 2025; Yuan et al., 2023; Iloanusi and Chun, 2024). The LLMs driving much of current AI developments are known to encode many societal biases, including on race (Sap et al., 2019; Hofmann et al., 2024), gender (Ding et al., 2025), language (Alhanai et al., 2025), and culture (Santy et al., 2023). While these challenges and biases are known, there is work needed to determine if and how AI can be designed for LEP communities while taking into consideration current model limitations and existing healthcare support systems.
@@ 2.3. Considerations for Health Technologies in Marginalized Communities
While digital health technologies generally expand access to care (Raney et al., 2017; Al-Shorbaji, 2022), they are often underutilized by LEP patients (Nouri et al., 2020; Gordon et al., 2025; Hsueh et al., 2021), thus exacerbating existing inequalities (Yao et al., 2022; Linggonegoro et al., 2021). For example, during and after the COVID-19 pandemic, LEP patients utilized video visits and online patient portals at a lower rate than English-proficient patients (Rodriguez et al., 2021).
Several structural factors impede health technology access for LEP individuals. Limited reading, digital, and health literacy can restrict meaningful engagement with digital tools such as patient portals (Lazaro et al., 2024; Smith and Magnani, 2019). LEP populations face generally lower rates of access to technology (Higashi et al., 2025), but even when access is available, they report distrust in digital services and difficulty understanding privacy policies (Tan-McGrory et al., 2022), fueling their hesitation in adopting technology that solicits their personal information. Recent legal actions and policies in the U.S. have compounded these concerns (Page and Polk, 2017; Fleming et al., 2019), underscoring the need to prioritize privacy and trust in equitable technology design. There are, however, community-based health service approaches that have proven to be accessible and productive for LEP individuals. SMS programs, WhatsApp groups, and community health worker (CHW)-linked tools are used by Latino immigrant communities and have been shown to support self-management and information access (Aibo et al., 2024; Anaya et al., 2021).
In consideration of the barriers, benefits, and risks for technology to support LEP individuals in healthcare, it is important to understand how these technologies can be implemented effectively and safely. There is little recent work that explores all of these factors together for LEP populations in particular. Accordingly, we report our findings around the barriers and experiences that LEP individuals may face when receiving healthcare, and in what ways AI technology can be designed to facilitate communication with their healthcare teams.
@@ 3.1. Study Context
In this study, we interviewed patient navigators who support Spanish-speaking LEP individuals when interacting with healthcare systems in the U.S. We focused on patient navigators, rather than patients themselves, because as we reflected on the tensions involved in conducting research with a marginalized community (in this case, LEP individuals seeking care), we were concerned that starting with LEP individuals might risk perpetrating further harm to the community by taking data and information and providing little in return (Liang et al., 2021). Speaking directly with patients at the outset could also risk misrepresenting or excluding key aspects of their experiences if we failed to frame the right questions or if patients were hesitant to disclose sensitive information, especially since the lead researcher in the study was not themselves a part of this community (Section 3.7). In contrast, patient navigators routinely accompany many patients through the care-seeking process, providing crucial perspectives that surface both common experiences and specific instances that reveal how patients engage with healthcare and technology. Talking to navigators allowed us to gain an understanding of the culture and background of LEP individuals, identify potential avenues for giving back to the community (e.g., presenting at community events, conducting follow-up design workshops), and build relationships with these communities for further research into technologies that are sensitive to the needs and lived experiences of LEP individuals. We recognize that navigators themselves may also be part of LEP communities or other marginalized groups. Accordingly, we worked closely with three navigators to provide feedback on pilot studies and assess what value such a study might bring to navigators. We discuss future work building on our findings to bring benefits of new AI health tools to both LEP individuals and patient navigators (Section 5).
Initial feedback from navigators was that LEP individuals vary widely in their experience of seeking care due to their range of languages, backgrounds, and cultures. In response to this feedback, we focused on Spanish-speaking LEP individuals for two reasons: Spanish is by far the most commonly spoken language at home after English in the U.S. (U.S. Census Bureau, 2022), and the majority of organizations we worked with focused primarily on Spanish speakers, giving us greater access to navigators and community resources with this population. However, as we will further discuss in our findings (Section 4.1), even with a shared language, individuals had varying experiences and identities (e.g., be immigrants from various countries or U.S.-born). In the rest of this paper, we may refer to Spanish-speaking LEP individuals as just LEP individuals, for simplicity.

### v2

@@ Abstract.
Limited English proficiency (LEP) patients in the U.S. face systemic barriers to healthcare beyond language and interpreter access, encompassing procedural and institutional constraints. AI advances may support communication and care through on-demand translation and visit preparation, but also risk exacerbating existing inequalities. We conducted storyboard-driven interviews with 14 patient navigators to explore how AI could shape care experiences for Spanish-speaking LEP individuals. We identified tensions around linguistic and cultural misunderstandings, privacy concerns, and opportunities and risks for AI to augment care workflows. Participants highlighted structural factors that can undermine trust in AI systems, including sensitive information disclosure, unstable technology access, and low literacy. While AI tools can potentially alleviate social barriers and institutional constraints, there are risks of misinformation and reducing human-to-human interactions. Our findings contribute AI design considerations that support LEP patients and care teams via rapport-building, educational and language support, and minimizing disruptions to existing practices. [^1] Supplementary materials available at https://github.com/mhuang412/designing-beyond-language
@@ 1. Introduction
Patient-provider communication is strongly linked to health outcomes and quality of care (Arora, 2003; Street Jr et al., 2009; Ong et al., 2000). Effective communication builds therapeutic relationships, satisfaction, trust, and long-term health outcomes (Street Jr et al., 2009; Ha and Longnecker, 2010), while poor communication contributes to errors, compromised patient safety, and inefficient use of resources (Vermeir et al., 2015; Tiwary et al., 2019). Language barriers in particular can exacerbate these challenges. In the U.S., over 25 million individuals aged 5 years and older report speaking English less than “very well” (U.S. Census Bureau, 2022). These individuals with limited English proficiency (LEP) are more likely to experience misdiagnoses, medication complications, poorer adherence to treatment plans, and decreased comprehension of diagnoses (Crane, 1997; Gandhi et al., 2000; Flores et al., 2006; Tiwary et al., 2019). Consequently, patient-provider communication, relationship-building, and shared decision-making break down (Paredes et al., 2018; Flores et al., 2006), reducing both patient and provider satisfaction (Al Shamsi et al., 2020).
@@
Digital health technologies, such as telehealth, online patient portals, and mobile health applications, can facilitate asynchronous communication and on-demand information access, which in principle could improve the quality and convenience of healthcare by enabling continuous and accessible support (Fatehi et al., 2020; Madanian et al., 2023; Ezeamii et al., 2024). However, LEP populations underutilize these resources due to persistent implementation barriers such as low digital and health literacy, unreliable Internet access, and ethical and privacy concerns (Ahmed et al., 2025; Khan et al., 2025; Madanian et al., 2023; Nouri et al., 2020; Payvandi et al., 2022). Even technologies designed specifically to accommodate LEP individuals, such as multilingual patient portals, are often poorly integrated or have limited functionalities, thus discouraging use (Anthony et al., 2024).
@@
Emerging AI-powered health technologies have the potential to provide personalized and timely medical support and translation (Johnson et al., 2021; Genovese et al., 2024; Miller and Brown, 2018), enabling LEP patients to engage with the healthcare system more effectively. For example, on-demand AI translation can facilitate patient-provider conversations (Mehandru et al., 2023), and new large language model (LLM)-powered chatbots can summarize and simplify complex medical information (Guo et al., 2021) or help patients prepare for visits (Liu et al., 2025; Samiee et al., 2025). However, most AI-powered interventions target digitally literate, English-speaking patients (Jin et al., 2024; Naamati-Schneider and Alt, 2024), and the underlying models generally struggle to adapt to marginalized cultural (Santy et al., 2023) and language backgrounds (Basoah et al., 2025; Sap et al., 2019). Furthermore, little is known about how AI interventions should be designed and governed when language barriers intersect with cultural norms, privacy fears, and low literacy, especially in contexts where human intermediaries (e.g., patient navigators) already play a critical role. Therefore, the rapid deployment of such AI health interventions risks replicating or further exacerbating existing disparities (Ahmed et al., 2025).
@@ 2. RQ2:
To address the above RQs, we conducted 14 storyboard-centered interviews with individuals who support Spanish-speaking LEP individuals receiving care in the U.S. These individuals, whom we collectively refer to as patient navigators, assisted patients in various ways, including accompanying them to provider visits, translating, teaching digital skills, and bridging resources. Their role as mediators between patients and providers lends them insights into the broader contexts and relational dynamics that shape LEP patient experiences. Consequently, they build on their experiences to reason about the feasibility and risks of developing AI technologies to facilitate equitable communication and care.
From these interviews, we identified key communication and systemic challenges that Spanish-speaking LEP patients face, such as low literacy and privacy concerns. These challenges also highlighted possible risks and opportunities for using AI, such as reducing social barriers to care while potentially diminishing existing human support. Our findings underscore the importance of adapting to cultural and linguistic nuances, overcoming various types of literacy barriers, and respecting user privacy to foster trust. These findings inform practical guidelines for developing AI tools that support LEP populations. Our contributions are summarized as follows:
@@ 2.1. Mediating Clinical Communication with LEP Individuals
Effective healthcare requires two-way communication in which providers contribute medical expertise and patients share their needs and preferences (Iroegbu et al., 2025; Williams et al., 2017). Strong interpersonal communication has been shown to influence patient outcomes as much as, if not more than, clinical instruction (Bartlett et al., 1984), leading to improved adherence to treatment plans (Street Jr et al., 2009), understanding of diagnoses (Kennedy MBA et al., 2014), and satisfaction (Ruben et al., 2020). Conversely, impeded communication is linked to more errors (Tiwary et al., 2019), worsened decision-making (Paredes et al., 2018), and lower rates of physician visits and preventive services (Jacobs et al., 2006; Woloshin et al., 1997; Fiscella et al., 2002). For LEP individuals, these challenges are compounded by linguistic (Schouten et al., 2020), cultural (Lazaro et al., 2024), intersectional (Parthab Taylor et al., 2013), and procedural (Frank, 2000) barriers. Therefore, there is growing interest in improving health outcomes by studying communication strategies that integrate patients’ lived experiences and values (Seo et al., 2021; Chen et al., 2025).
Research shows that language-concordant care, whether via direct communication  (Molina and Kasper, 2019) or interpretation  (Dunlap et al., 2015; Eskes et al., 2013), leads to fewer communication errors (Karliner et al., 2007) and higher satisfaction and perceived quality of care (Diamond et al., 2019) for LEP individuals. Despite these benefits, interpretation services remain underutilized due to barriers such as limited availability, cost, time constraints, and policies (Khoong and Fernandez, 2021; Jaeger et al., 2019; Jacobs et al., 2004). Even when interpretation is available, institutional policies can constrain interpreters to verbatim translation rather than clarification or advocacy, leaving unresolved misunderstandings and limiting relationship-building—conditions that shape what AI can realistically improve. Beyond linguistic access, broader sociocultural factors complicate LEP communication. Patients may self-censor due to stigma, fear of judgment, or mistrust of institutions (Levy et al., 2018; El-Krab et al., 2023). These dynamics highlight the need for cultural sensitivity, empathy, and attention to the broader contexts of patients’ lives. Patient navigators and community health workers (CHWs) function as critical infrastructure for marginalized communities, supporting interpretation, care coordination, documentation, and advocacy (Khatri et al., 2024). Prior work highlights how these roles reduce access barriers but also create invisible labor and emotional burdens (Verdezoto et al., 2021). Thus, understanding how AI might support or destabilize this ecosystem is necessary for equitable deployment.
@@ 2.2. AI Tools in Healthcare
Recent AI advancements have potential to support healthcare systems by improving clinical decision-making and workflows, resource allocations, medical imaging, patient monitoring, medical translations, and patient-provider relationships (Maleki Varnosfaderani and Forouzanfar, 2024; Shaheen, 2021; Barwise et al., 2024; Koski and Murphy, 2021). However, prior work has also identified key challenges and risks when implementing AI for healthcare, including data privacy, lack of transparency in decision-making, misinformation, bias mitigation, and cultural sensitivity (Maleki Varnosfaderani and Forouzanfar, 2024; Germani et al., 2024; Hussain, 2024; Capraro et al., 2024; Chustecki, 2024; Shi et al., 2025; Yoo et al., 2025).
AI has potential to support marginalized communities in receiving care by overcoming language barriers, providing personalized health information, and offering support in rural or low-resource areas (Genovese et al., 2024; Kaur and Mekled, 2025). Barwise et al. (2024) explored patients’ perception of using predictive models for interpreter prioritization and Mehandru et al. (2023) used physician-in-the-loop methods to improve reliance of machine translation in settings where no human translators were available. Other work has also explored using chatbots for pre-clinical consultations (Liu et al., 2025; Samiee et al., 2025) and helping patients read physician notes (Kambhamettu et al., 2025), discharge instructions (Zaretsky et al., 2024), and medical information (Guo et al., 2021). While AI presents many exciting possibilities to support patient care, the majority of such interventions are focused on English or translating into English, posing issues for low-resource languages such as reinforcing biases and overlooking cultural and infrastructural challenges that may cause AI to exacerbate rather than alleviate inequities in underserved populations (Wasi, 2025; Yuan et al., 2023; Iloanusi and Chun, 2024). The LLMs driving much of current AI developments are known to encode biases on race (Sap et al., 2019; Hofmann et al., 2024), gender (Ding et al., 2025), language (Alhanai et al., 2025), and culture (Santy et al., 2023). While these challenges and biases are known, there is work needed to determine if and how AI can be designed for LEP communities while taking into consideration current model limitations and existing healthcare support systems.
@@ 2.3. Considerations for Health Technologies in Marginalized Communities
While digital health technologies generally expand access to care (Raney et al., 2017; Al-Shorbaji, 2022; Suh et al., 2024), they are often underutilized by LEP patients (Nouri et al., 2020; Gordon et al., 2025; Hsueh et al., 2021), thus exacerbating existing inequalities (Yao et al., 2022; Linggonegoro et al., 2021). For example, during and after the COVID-19 pandemic, LEP patients utilized video visits and online patient portals at a lower rate than English-proficient patients (Rodriguez et al., 2021).
Several structural factors impede health technology access and utilization for LEP individuals. Limited reading, digital, and health literacy can restrict meaningful engagement with digital tools (Lazaro et al., 2024; Smith and Magnani, 2019), yet digital health literacy is key to realizing equitable advancements in healthcare (Bywall et al., 2025). Prior work in HCI also highlights the need for AI technologies to be culturally sensitive (Deva et al., 2025; Vigil-Hayes et al., 2021) in order to be more trustworthy and accessible to marginalized users. Furthermore, LEP populations face generally lower rates of access to technology (Higashi et al., 2025), but even when access is available, they report distrust in digital services and difficulty understanding privacy policies (Tan-McGrory et al., 2022), fueling their hesitation in adopting technology that solicits their personal information. Recent legal actions and policies in the U.S. have compounded these concerns (Page and Polk, 2017; Fleming et al., 2019), underscoring the need to prioritize privacy and trust in equitable technology design. There are, however, community-based health service approaches that have proven to be accessible and productive for LEP individuals. SMS programs, WhatsApp groups, and CHW-linked tools used by Latino immigrant communities have been shown to support self-management and information access (Aibo et al., 2024; Anaya et al., 2021).
In consideration of the barriers, benefits, and risks for technology to support LEP individuals in healthcare, it is important to understand how these technologies can be implemented equitably and safely (Parker et al., 2025). There is little recent work that explores all of these factors together for LEP populations in particular. Accordingly, we report our findings around the barriers and experiences that LEP individuals may face when receiving healthcare, and in what ways AI technology can be designed to facilitate communication with their healthcare teams.
@@ 3.1. Study Context
In this study, we interviewed patient navigators who support Spanish-speaking LEP individuals when interacting with healthcare systems in the U.S. We focused on patient navigators, rather than patients themselves, because as we reflected on the tensions involved in conducting research with a marginalized community (in this case, LEP individuals seeking care), we were concerned that starting with LEP individuals might risk perpetrating further harm to the community by taking data and information and providing little in return (Liang et al., 2021). Speaking directly with patients at the outset could also risk misrepresenting or excluding key aspects of their experiences if we failed to frame the right questions or if patients were hesitant to disclose sensitive information, especially since the lead researcher in the study was not themselves a part of this community (Section 3.7). In contrast, patient navigators routinely accompany many patients through the care-seeking process, providing crucial perspectives that surface both recurring experiences and specific incidents that reveal how patients engage with healthcare and technology. Talking to navigators allowed us to gain an understanding of the culture and background of LEP individuals, identify potential avenues for giving back to the community (e.g., presenting at community events, conducting follow-up design workshops), and build relationships with these communities for further research into technologies that are sensitive to the needs and lived experiences of LEP individuals. We recognize that navigators themselves may also be part of LEP communities or other marginalized groups. Accordingly, we worked closely with three navigators to provide feedback on pilot studies and assess what value such a study might bring to navigators. We discuss future work building on our findings to bring benefits of new AI health tools to both LEP individuals and patient navigators (Section 5).
Initial feedback from navigators was that LEP individuals vary widely in their care experiences due to the range of languages, backgrounds, and cultures. In response to this feedback, we focused on Spanish-speaking LEP individuals for two reasons: Spanish is by far the most commonly spoken language at home after English in the U.S. (U.S. Census Bureau, 2022), and the majority of organizations we worked with focused primarily on Spanish speakers, giving us greater access to navigators and community resources with this population. However, as we will further discuss in our findings (Section 4.1), even with a shared language, individuals had varying experiences and identities (e.g., be immigrants from various countries or U.S.-born). In the rest of this paper, we refer to Spanish-speaking LEP individuals as just LEP individuals, for simplicity.